Select one card from a standard deck. What is the probability of selecting a four, seven, or a heart?
Answer by checkley77(12844)

hearts=13
4's=3 (4 of hearts already included)
7's=3 (7 of hearts already included)
succes/total=(13+3+3)/52=19/52 or.3654 or 36.54% probability of success.